Suppose that a random variable X is normally distributed with a mean of 200 and a variance of 

625.

Required: 

a) What proportion of X-values lies between 180 and 195? (7) 

b) Below what value do 33 per cent of X-values lie? (6)

c) What proportion of X-values are more than 190? (6)

d) What proportion of X-values are less than 195? (6)

~END~


1
Expert's answer
2021-04-14T12:51:27-0400

a)

"z=\frac{X-\mu}{\sigma}"

"\sigma=\sqrt{625}=25"

"z_1=\frac{180-200}{25}=-0.8"

"z_2=\frac{195-200}{25}=-0.2"

"P(180<X<195)=P(z<-0.2)-P(z<-0.8)=0.4207-0.2119=0.2088"


b)

"P(z<-0.44)=0.33=33\%"

"z=\frac{X-200}{25}=-0.44"

"X=-0.44\cdot25+200=189"


c)

"z=\frac{190-200}{25}=-0.4"

"P(X>190)=1-P(z<-0.4)=1-0.3446=0.6554"


d)

"z=\frac{195-200}{25}=-0.2"

"P(X<195)=P(z<-0.2)=0.4207"